OpenStudy (anonymous):

Find the dot product, a · b. a = 6i + 5j, b = -5i + 4j

OpenStudy (freckles):

multiply the scalars of the i components multiple the scalars of the j components add both of those products to find the dot product of a and b

OpenStudy (johnweldon1993):

Right, as said above \[\large \vec{a} = <a_i, a_j>\] \[\large \vec{b} = <b_i, b_j>\] \[\large a \cdot b = (a_i \times b_i + a_j \times b_j)\]